# Quantifying Interpretability in CLIP Models with Concept Consistency

### v1

@@ 4.3.2 Models
5 Results and Discussion
5.1 Concept Consistency Scores (CCS) for CLIP models.
@@ 5 Results and Discussion
5.2 Interpretable CLIP Models: The Role of CCS.
In the previous section, we analyzed the distribution of CCS scores across various CLIP models. This naturally raises an important question: How does CCS provide deeper insights into the functional role of individual attention heads in influencing downstream tasks? To address this, we conduct a soft-pruning analysis targeting attention heads with extreme CCS scores — specifically, heads with high CCS scores (i.e., CCS = 5) and low CCS scores (CCS $\leq$ 1). Soft-pruning is implemented by zeroing out the attention weights of selected heads, thereby nullifying their contribution to the model’s predictions without altering the overall model architecture.
Tables  4 and  5 report zero-shot classification performance on CIFAR-10, CIFAR-100, FOOD-101, Country-211, and Oxford-Pets, under three conditions: no pruning (original), pruning of high-CCS heads, and pruning of low-CCS heads. We observe that pruning heads with high CCS scores consistently leads to a significant degradation in performance across all datasets, whereas pruning low CCS heads has a much smaller impact. This clear difference in sensitivity demonstrates that CCS effectively identifies heads that capture critical, concept-aligned information essential for model predictions. Thus, CCS serves as a reliable indicator of the most functionally important heads in CLIP models, offering a concrete step toward interpreting the internal mechanisms of these large vision-language models.
Interestingly, we also observe variations in sensitivity across model architectures. The ViT-B-16 models experience the most substantial performance drops when high CCS heads are pruned, suggesting a higher dependency on a smaller set of concept-specialized heads. In contrast, ViT-L-14 models exhibit the least drop, indicating a more distributed representation. Among smaller models, OpenAI-trained models show a larger performance degradation compared to OpenCLIP counterparts when high CCS heads are removed, whereas for larger models like ViT-L-14, OpenCLIP models exhibit a slightly higher drop than OpenAI models. These trends further highlight how CCS can reveal architectural and training-specific differences in how CLIP models encode and utilize concepts for downstream tasks.
5.3 High CCS vs random heads pruning
To rigorously evaluate the effectiveness of the Concept Consistency Score (CCS) as a measure of interpretability in CLIP models, we compare the impact of pruning heads with high CCS scores against pruning an equal number of randomly selected heads. While earlier results demonstrated that pruning high CCS heads significantly degrades model performance, a critical question remains: Are these high-CCS heads genuinely more important than other heads?
To investigate this, we conduct controlled experiments where we randomly prune the same number of attention heads and analyze the corresponding performance drop. The results of these experiments are presented in Figure 2. As shown in the figure, pruning high CCS heads consistently leads to a substantially larger decrease in zero-shot performance compared to random pruning, across different datasets and model variants. These findings empirically validate that CCS effectively identifies heads that are essential for the model’s decision-making process, thereby offering a principled mechanism for interpreting the internal workings of CLIP models. Unlike random pruning, which affects heads without regard to their learned properties, CCS-guided pruning systematically targets heads that encode critical concepts, revealing their functional role in model predictions.
Moreover, we observe that larger CLIP models exhibit smaller differences between high CCS and random pruning impacts compared to smaller models, suggesting that larger architectures may possess more redundancy or distributed representations, making them more resilient to head pruning. Overall, these results establish CCS as a reliable and interpretable metric for identifying concept-relevant heads, contributing to a deeper understanding of how CLIP models organize and utilize conceptual knowledge across their layers.
5.4 High CCS heads are crucial for out-of-domain (OOD) detection
While our earlier experiments primarily focused on in-domain datasets such as CIFAR-10 and CIFAR-100 to validate the Concept Consistency Score (CCS), understanding model behavior under out-of-domain (OOD) conditions is a critical step toward evaluating models’ robustness and spurious correlations. Table  6 demonstrates the results on ImageNet-A and ImageNet-R datasets respectively. From the table, we observe that pruning heads with high CCS scores leads to a substantial degradation in model performance, underscoring the critical role these heads play in the model’s decision-making process. Notably, the ViT-B-16-OpenAI model exhibits the most pronounced drop in performance upon pruning high CCS heads, suggesting that this model relies heavily on a smaller set of concept-specific heads for robust feature representation consistent with the observations previously. These results demonstrate that CCS is a powerful metric for identifying attention heads that encode essential, generalizable concepts in CLIP models while avoiding spurious correlations.
(a) MSRVTT
(b) MSVD
(c) DIDEMO
5.5 Impact of High CCS heads on corresponding tasks.
To further investigate the functional role of high CCS heads, we perform concept-specific pruning experiments. Specifically, we prune heads with high CCS scores for a target concept (e.g., locations) and evaluate the model on tasks aligned with that concept (e.g., location classification). To contrast this, we also prune heads associated with unrelated concepts (e.g., animals) and compare the impact on task performance. The motivation behind this analysis is to validate whether CCS-identified heads are not only aligned with specific concepts but also crucial for performing tasks involving those concepts. Figure  3 and  4 shows the zero-shot results on Country-211 and CIFAR10-objects datasets. These datasets primarily test varied classes of countries and objects respectively.
5.5.1 High CCS heads are crucial for concept-aligned tasks.
Pruning heads with high CCS scores for a target concept (e.g., locations) leads to a significant drop in task performance when the task is aligned with that concept. For example, in ViT-B-16 (openai), location classification accuracy drops from 22.81% to 14.09% when location heads are pruned. This clearly shows that CCS accurately identifies heads encoding concept-relevant information, making them essential for corresponding tasks.
5.5.2 Minimal Impact of Unrelated Head Pruning.
When pruning heads associated with concepts unrelated to the task (e.g., pruning animal heads for location classification), the model’s performance is largely unaffected, demonstrating the concept specificity of high CCS heads.
5.5.3 Objects heads are the most critical for general classification.
In general classification tasks, pruning object-related heads consistently reduces accuracy more than pruning location or color heads. For example, In ViT-B-32 (openai), accuracy drops from 87.6% to 86.02% with object head pruning, while location and color head pruning show smaller drops (87.02% and 87.22%, respectively). This indicates that object heads are more critical to overall recognition tasks, aligning with the importance of object features in vision models.
5.5.4 Larger models show robustness to pruning.
Larger models like ViT-L-14 exhibit smaller performance drops when pruning concept-specific heads, suggesting more distributed and redundant representations. For example, In ViT-L-14 (openai), object head pruning reduces accuracy marginally from 92.1% to 91.25%, and location/color head pruning has negligible effect.
These results validate the effectiveness of CCS as an interpretability tool, demonstrating that heads with high CCS scores: (i). Encode semantically aligned and task-relevant information, (ii) Are critical for model decision-making in tasks involving those concepts, and (iii). Offer targeted insights into how concepts are represented inside CLIP models. Additionally, larger models’ robustness to pruning highlights differences in representation capacity and redundancy across architectures.
5.6 Impact of CCS pruning on zero-shot video retrieval.
6 Conclusion
In this work, we proposed Concept Consistency Score (CCS), a novel interpretability metric that quantifies how consistently individual attention heads in CLIP-like models align with semantically meaningful concepts. Through extensive soft-pruning experiments, we demonstrated that heads with high CCS are essential for maintaining model performance, as their removal leads to substantial performance drops compared to pruning random or low CCS heads. Our findings further highlight that high CCS heads are not only critical for standard vision-language tasks but also play a central role in out-of-domain detection and concept-specific reasoning. Moreover, experiments on video retrieval tasks reveal that high CCS heads are crucial for capturing temporal and cross-modal relationships, underscoring their broad utility in multimodal understanding. Our study shows how CCS is a powerful interpretability metric for identifying key layers and heads in CLIP-like models.

### v2

@@ 4.2 High CCS vs random heads pruning
(a) MSRVTT
(b) MSVD
(c) DIDEMO
4.3 High CCS heads are crucial for out-of-domain (OOD) detection
While our earlier experiments primarily focused on in-domain datasets such as CIFAR-10 and CIFAR-100 to validate the Concept Consistency Score (CCS), understanding model behavior under out-of-domain (OOD) conditions is a critical step toward evaluating models’ robustness. Table  4 demonstrates the results on ImageNet-A and ImageNet-R datasets respectively. From the table, we observe that pruning heads with high CCS scores leads to a substantial degradation in model performance, underscoring the critical role these heads play in the model’s decision-making process. Notably, the ViT-B-16-OpenAI model exhibits the most pronounced drop in performance upon pruning high CCS heads, suggesting that this model relies heavily on a smaller set of concept-specific heads for robust feature representation consistent with the observations previously. These results demonstrate that CCS is a powerful metric for identifying attention heads that encode essential, generalizable concepts in CLIP models.
4.4 High CCS heads are crucial for concept-specific tasks.
To investigate the functional role of high Concept Consistency Score (CCS) heads, we conduct concept-specific pruning experiments. In these experiments, we prune heads with high CCS scores corresponding to a target concept (e.g., locations) and evaluate the model’s performance on tasks aligned with that concept, such as location classification. In contrast, we also prune heads associated with unrelated concepts (e.g., animals) and assess the resulting impact on task performance. Our results indicate that pruning high CCS heads leads to a significant drop in task performance, validating that these heads encode essential concept-relevant information. For instance, in the ViT-B-16 model, pruning location heads results in a substantial decrease in location classification accuracy from 22.81% to 14.09%, as shown in Figure  3. Conversely, pruning heads corresponding to unrelated concepts has little effect on performance, demonstrating the concept-specific nature of high CCS heads, as illustrated in Figure  4.
In more general classification tasks, object-related heads consistently exhibit a greater impact on performance than location or color heads. For example, in the ViT-B-32 model, pruning object-related heads leads to a more noticeable accuracy drop (from 87.6% to 86.02%) compared to pruning location or color heads, which result in smaller reductions (87.02% and 87.22%, respectively). This underscores the greater importance of object-related features in vision tasks. Larger models, such as ViT-L-14, demonstrate a more robust performance to pruning, with smaller accuracy drops when pruning concept-specific heads, suggesting that these models employ more distributed and redundant representations. For instance, pruning object-related heads in ViT-L-14 reduces accuracy only marginally, from 92.1% to 91.25%, with negligible effects from pruning location and color heads. These results not only confirm the effectiveness of CCS as an interpretability tool but also show that high CCS heads are critical for concept-aligned tasks and provide significant insights into how concepts are represented within CLIP-like models.
4.5 Impact of CCS pruning on zero-shot video retrieval.
Occupations
biologist, composer, economist, mathematician, model, poet, reporter, zoologist, artist, coach, athlete, audiologist, judge, musician, therapist, banker, ceo, consultant, prisoner, assistant, boxer, commander, librarian, nutritionist, realtor, supervisor, architect, priest, guard, magician, producer, teacher, lawyer, paramedic, researcher, physicist, pediatrician, surveyor, laborer, statistician, dietitian, sailor, tailor, attorney, army, manager, baker, recruiter, clerk, entrepreneur, sheriff, policeman, businessperson, chief, scientist, carpenter, florist, optician, salesperson, umpire, painter, guitarist, broker, pensioner, soldier, astronaut, dj, driver, engineer, cleaner, cook, housekeeper, swimmer, janitor, pilot, mover, handyman, firefighter, accountant, physician, farmer, bricklayer, photographer, surgeon, dentist, pianist, hairdresser, receptionist, waiter, butcher, videographer, cashier, technician, chemist, blacksmith, dancer, doctor, nurse, mechanic, chef, plumber, bartender, pharmacist, electrician
Table 5: Full list of occupations used for evaluating biases on FairFace and SocialCounterFactuals datasets.
Prompt Example
A <occupation> A biologist
A photo of <occupation> A photo of biologist
A picture of <occupation> A picture of biologist
An image of <occupation> An image of biologist
Table 6: Prompts used for measuring biases on FairFace and SocialCounterFactuals datasets.
Model Race Gender
Original High CCS Original High CCS
ViT-B-32-OpenAI 61.75 60.47 41.24 41.11
ViT-B-32-datacomp 49.2 48.32 21.35 20.97
ViT-B-16-OpenAI 64.61 55.55 40.19 38.21
ViT-B-16-LAION 59.21 56.72 43.55 43.11
ViT-L-14-OpenAI 59.28 59.75 34.7 32.23
ViT-L-14-LAION 61.92 55.59 43.02 39.17
Table 7: Comparison of original and high-CCS pruning on FairFace dataset for race and gender.. We used MaxSkew (K=900) as the metric.
Model Race Gender
Original High CCS Original High CCS
ViT-B-32-OpenAI 3.65 2.43 4.05 1.22
ViT-B-16-OpenAI 2.43 1.22 0.81 2.03
ViT-L-14-OpenAI 2.03 0.81 2.42 1.62
Table 8: Comparison of original and high-CCS soft-pruning on SocialCounterFactuals dataset for race and gender.. We used MaxSkew (K=12 for race, K=4 for gender) as the metric.
4.6 CLIP’s high-CCS heads encode features that drive social biases.
Previously, we established that high-CCS heads in CLIP models are crucial for image and video tasks and pruning them leads to significant drop in performance. Now, we investigate if these high CCS heads learn spurious features leading to social biases. For this, we perform soft pruning experiment on FairFace and SocialCounterFactuals datasets. Here given neutral text prompts of 104 occupations as show in tables  5 and  6, we measure MaxSkew across race and gender in the datasets. Tables  7 and  8 show the results on FairFace and SocialCounterFactuals datasets respectively.
On the FairFace dataset, pruning high-CCS heads consistently reduces the MaxSkew values for both race and gender across all models. For example, in the ViT-B-16-OpenAI model, pruning high-CCS heads drops the race MaxSkew from 64.61 to 55.55 and the gender MaxSkew from 40.19 to 38.21. Similar reductions are observed across all ViT-B and ViT-L variants. These drops, although modest in some cases, indicate a consistent trend: high-CCS heads are contributing disproportionately to skewed model predictions. The effect is even more evident on the SocialCounterfactuals dataset, where MaxSkew values drop substantially upon pruning high-CCS heads. For instance, in ViT-B-32-OpenAI, the gender MaxSkew falls from 4.05 to 1.22, and race MaxSkew from 3.65 to 2.43. Similar reductions occur for other ViT variants, with some pruned models showing more than 50% decrease in bias.
These results reveal a fundamental paradox at the heart of CLIP models: high-CCS heads, while critical for strong performance in tasks such as classification, retrieval, and concept alignment, are also the primary contributors to social bias. This paradox emerges from CLIP’s contrastive learning objective, which optimizes alignment between images and their paired text across large, uncurated datasets. In doing so, the model often absorbs and amplifies spurious correlations between visual features and demographic attributes. High-CCS heads, by virtue of their consistent focus on semantically aligned regions, become particularly susceptible to reinforcing these correlations. Pruning these heads leads to a notable reduction in model bias, as shown in our experiments, but also comes at the cost of reduced performance—a clear trade-off between fairness and utility. This performance-bias paradox underscores the complex role of high-CCS heads: they are both enablers of semantic understanding and carriers of learned stereotypes. The CCS metric, in this context, provides a valuable lens for navigating this tension. It not only aids in interpreting model behavior but also offers a lightweight intervention—soft-pruning—that mitigates bias without requiring expensive fine-tuning.
5 Conclusion
In this work, we proposed Concept Consistency Score (CCS), a novel interpretability metric that quantifies how consistently individual attention heads in CLIP-like models align with semantically meaningful concepts. Through extensive soft-pruning experiments, we demonstrated that heads with high CCS are essential for maintaining model performance, as their removal leads to substantial performance drops compared to pruning random or low CCS heads. Our findings further highlight that high CCS heads are not only critical for standard vision-language tasks but also play a central role in out-of-domain detection and concept-specific reasoning. Moreover, experiments on video retrieval tasks reveal that high CCS heads are crucial for capturing temporal and cross-modal relationships, underscoring their broad utility in multimodal understanding. In addition, we demonstrated that high-CCS heads learn spurious correlations leading to social biases and pruning them mitigates that harmful behaviour without the need for further finetuning. Thus, CCS provides an wholistic view of interpretability proving the paradox performance vs social biases in CLIP.
Appendix A Concept Consistency Scores (CCS) for CLIP models.
@@ 5 Conclusion
ViT-B-32-OpenAI
L8.H11 (Descriptive), L9.H2 (Objects), L9.H3 (Descriptions), L10.H8 (Locations), L11.H1 (Objects), L11.H5 (Colors), L11.H7 (Objects), L11.H9 (Locations)
ViT-B-32-datacamp
L8.H1 (Objects), L8.H3 (Subjects), L8.H10 (Objects), L9.H3 (Subjects), L9.H10 (Objects), L10.H7 (Locations), L10.H11 (Objects), L11.H3 (Colors), L11.H4 (Colors), L11.H9 (Colors), L11.H10 (Objects)
ViT-B-16-OpenAI
L8.H5 (Visual), L8.H8 (Visual), L10.H5 (Subjects), L10.H7 (Settings), L11.H0 (Creative), L11.H3 (Settings), L11.H4 (Stylistic), L11.H6 (Locations), L11.H7 (Colors), L11.H11 (Animals)
ViT-B-16-LAION
L8.H6 (Descriptions), L8.H7 (Descriptions), L9.H0 (Themes), L9.H1 (Aesthetics), L9.H3 (Descriptive), L10.H5 (Artwork), L10.H10 (Locations), L11.H0 (Locations), L11.H2 (Descriptions), L11.H6 (Locations), L11.H7 (Objects), L11.H8 (Objects), L11.H10 (Colors)
ViT-L-14-OpenAI
L20.H2 (Locations), L20.H12 (Descriptions), L21.H0 (Locations), L21.H1 (Locations), L21.H8 (Expressions), L21.H13 (Locations), L21.H15 (Locations), L22.H1 (Objects), L22.H2 (Locations), L22.H5 (Locations), L22.H9 (Subjects), L22.H13 (Animals), L22.H15 (Locations), L23.H4 (Objects), L23.H10 (Locations), L23.H11 (Colors)
ViT-L-14-LAION
L20.H4 (Subjects), L20.H14 (Descriptions), L21.H0 (Colors), L21.H1 (Locations), L21.H5 (Descriptive), L21.H9 (Colors), L21.H11 (Locations), L22.H0 (Patterns), L22.H1 (Shapes), L22.H3 (Objects), L22.H5 (Visual), L22.H6 (Animals), L22.H8 (Letters), L22.H10 (Colors), L22.H12 (Landscapes), L22.H13 (Locations), L23.H4 (People), L23.H5 (Nature), L23.H6 (Locations), L23.H8 (Colors), L23.H9 (Descriptive)
Table 10: Full List of high-CCS heads of all CLIP models.
ViT-B-32-OpenAI
L8.H1 (Artistic), L8.H2 (Objects), L8.H6 (Photography), L8.H9 (Styles), L8.H10 (Perspective), L9.H1 (Subjects), L9.H11 (Settings), L10.H0 (Objects), L10.H3 (Locations), L10.H7 (Locations), L11.H6 (Descriptions), L11.H10 (Locations), L11.H11 (Locations)
ViT-B-32-datacamp
L8.H0 (Environments), L8.H7 (Creativity), L9.H6 (Colors), L10.H5 (Art), L10.H6 (Descriptions), L10.H8 (Locations), L10.H9 (Descriptions), L11.H2 (Subjects), L11.H8 (Qualities)
ViT-B-16-OpenAI
L8.H1 (Artistic), L8.H2 (Photography), L8.H4 (Styles), L8.H6 (Artwork), L8.H7 (Photography), L8.H9 (Light), L9.H4 (Photography), L9.H6 (Artforms), L9.H10 (Elements), L10.H3 (Locations), L10.H8 (Colors), L10.H9 (Artwork), L11.H5 (Objects), L11.H8 (Effects)
ViT-B-16-LAION
L8.H0 (Locations), L8.H8 (Text), L8.H9 (Photography), L9.H7 (Artistic), L9.H8 (Settings), L9.H11 (Descriptions), L10.H2 (Nature), L10.H3 (Location), L10.H7 (Expressions), L11.H3 (Settings), L11.H9 (Numbers), L11.H11 (Letters)
ViT-L-14-OpenAI
L20.H0 (Locations), L20.H3 (Locations), L20.H7 (Communication), L20.H8 (Vehicles), L20.H10 (Locations), L21.H4 (Photography), L21.H6 (People), L21.H10 (Locations), L22.H3 (Countries), L22.H12 (Professions), L23.H3 (Patterns), L23.H9 (Creativity), L23.H15 (Visual)
ViT-L-14-LAION
L20.H0 (Locations), L20.H1 (Locations), L20.H2 (Locations), L20.H8 (Locations), L20.H9 (Locations), L20.H11 (Aesthetics), L20.H15 (Descriptions), L21.H12 (Photography), L21.H14 (Locations), L22.H9 (Activities), L22.H14 (Colors), L22.H15 (Emotions), L23.H0 (Materials), L23.H3 (Settings)
Table 11: Full List of medium-CCS heads of all CLIP models.
ViT-B-32-OpenAI
L8.H5 (Patterns), L9.H9 (Ambiance), L11.H0 (Diverse), L11.H8 (Word)
ViT-B-32-datacamp
L8.H2 (Images), L8.H4 (Varied), L8.H9 (Varied), L9.H4 (Variety), L9.H5 (Professions), L11.H0 (Diverse), L11.H1 (Varied), L11.H11 (Settings)
ViT-B-16-OpenAI
L8.H0 (Diversity), L9.H3 (Locations), L10.H6 (Body parts), L11.H2 (Perspective)
ViT-B-16-LAION
L8.H4 (Variety), L8.H5 (Varied), L8.H10 (Diverse), L9.H2 (Textures), L10.H6 (Photography), L10.H8 (Traits)
ViT-L-14-OpenAI
L20.H1 (Diverse), L20.H4 (Diversity), L20.H6 (Items), L20.H15 (Diverse), L21.H2 (Diversity), L21.H3 (Diverse), L22.H0 (Occupations), L22.H4 (Settings), L22.H6 (Weather), L22.H14 (Items), L23.H5 (Diversity)
ViT-L-14-LAION
L20.H13 (Photography), L21.H6 (Professions), L23.H1 (Diverse)
Table 12: Full List of low-CCS heads of all CLIP models.